Help In the country of Economica, the total labor force consists of 10,000 workers 600 of these workers are unemployed, and thus the current unemployment rate is 6 0% This year, 200 workers will lose their jobs, 170 workers will return to the labor force after an absence. 100 workers will voluntarily leave their jobs, and 500 new workers will enter the labor force At the same time, 830 workers will find jobs, and 300 workers will retire or leave the labor force for other reasons Thus, at the end of the year, the labor force will consist ofworkers, and the unemployment rate will be Round your espo sest one de mal pace
Answer #1

Ans:

Labor force = employed workers + unemployed workers

Unemployed workers = 600

Employed workers = Labor force - Unemployed workers

= 10,000 - 600

= 9,400

Employed workers at the end of the year = 9,400 - 200 - 100 + 830

= 9,930

Unemployed workers at the end of the year = 600 + 200 + 500 - 830 - 300

= 170

1) Labor force at the end of the year = 9,930 + 170

= 10,100 workers

2) Unemployment rate = unemployed / labor force

= 170 / 10,100

= 1.7%